# Efficient Neural Common Neighbor for Temporal Graph Link Prediction

### v1

@@ Abstract
Temporal graphs are ubiquitous in real-world scenarios, such as social network, trade and transportation. Predicting dynamic links between nodes in a temporal graph is of vital importance. Traditional methods usually leverage the temporal neighborhood of interaction history to generate node embeddings first and then aggregate the source and target node embeddings to predict the link. However, such methods focus on learning individual node representations, but overlook the pairwise representation learning nature of link prediction and fail to capture the important pairwise features of links such as common neighbors (CN). Motivated by the success of Neural Common Neighbor (NCN) for static graph link prediction, we propose TNCN, a temporal version of NCN for link prediction in temporal graphs. TNCN dynamically updates a temporal neighbor dictionary for each node, and utilizes multi-hop common neighbors between the source and target node to learn a more effective pairwise representation. We validate our model on five large-scale real-world datasets from the Temporal Graph Benchmark (TGB), and find that it achieves new state-of-the-art performance on three of them. Additionally, TNCN demonstrates excellent scalability on large datasets, outperforming popular GNN baselines by up to 6.4 times in speed. Our code is available at https://github.com/GraphPKU/TNCN.
@@ 1 Introduction
Temporal graphs are increasingly utilized in contemporary real-world applications. Complex systems such as social networks (Yang et al., 2017; Min et al., 2021; Nguyen et al., 2017), trade and transaction networks (Zhang et al., 2018; Yan et al., 2021), and recommendation systems (Wu et al., 2022; Yin et al., 2019) are prime examples. These systems evolve dynamically over time, exhibiting different characteristics at various temporal moments. Recently, there has been a marked increase in the representation learning tasks on temporal graphs in the graph research community. At the same time, a prominent graph learning tool, Graph Neural Network (GNN) (Scarselli et al., 2008), has been developed to model node, link, and graph tasks. GNNs generally learn node embeddings by iteratively aggregating embeddings from neighboring nodes. They have demonstrated exceptional performance in graph representation learning, achieving state-of-the-art results across numerous tasks.
There remains a significant gap between static graphs and the increasingly prevalent temporal graphs. Temporal graphs incorporate discrete or continuous timestamps attached to edges, providing a more precise depiction of the graph evolution process. Meanwhile, many methodologies for static graphs are not applicable due to the additional constraints imposed by timestamps on causality. One can only utilize nodes and edges that precede a given time, implicitly resulting in numerous graph instances to analyze. Building on the success of time sequence modeling, Kumar et al. (2019); Trivedi et al. (2019); Rossi et al. (2020) propose memory-based temporal graph networks aimed at learning both short- and long-term dependencies. These methods, particularly Temporal Graph Networks (TGN), have achieved notable success in temporal node classification. Additionally, Transformer-based models, such as those proposed by Wang et al. (2021a); Xu et al. (2020), employ the multi-head attention mechanism to capture both cross and internal relationships within a temporal graph.
@@
However, these models may exhibit inherent flaws when addressing link prediction tasks in graphs. Such tasks require the model to predict the existence of a target link given specific prerequisite information. The aforementioned approaches primarily focus on node-wise representation learning, using the node embeddings of source and destination nodes to predict the existence of a target link. While straightforward, this method often fails to precisely capture the relationship between two nodes and other more complex graph structure features. For example, in Figure 1, the node-wise method might struggle to predict whether node $w$ prefers to interact with node $u$ or $v$. Generally speaking, these models are constrained by their focus on node-level representation, limiting their ability to capture broader contexts and higher-order patterns.
Considering these shortcomings, Zhang and Chen (2018) highlight the importance of pair-wise representation and propose labeling trick to mark the source and destination nodes when learning their pair-wise representation by GNN. Labeling trick (Zhang et al., 2021) is shown to greatly enhance GNN performance on static graph link prediction. Building on this, Wang et al. (2021c); Luo and Li (2022); Yu et al. (2023) develop models that leverage information from temporal surrounding nodes, successfully extending the pair-wise learning approach to dynamic graphs. They extract features using decreasing-timestamp random walks or joint neighborhood structures to generate multi-level embeddings for the center nodes, facilitating node classification and link prediction with pair-wise representations. However, these graph-based models often incur significant computational and memory costs due to extracting temporal neighborhoods and applying message passing on them for each node/link to predict, hindering their application in real-world, large-scale scenarios.
In general, graph-based representation learning methods can significantly enhance model capabilities, yet their high computational cost limits widespread application. For example, the Temporal Graph Benchmark (TGB) (Huang et al., 2023) contains many high-quality real-world temporal graphs. However, most graph-based models choose to only evaluate on a subset of small graphs due to their unaffordable cost of training and evaluation on large-scale datasets. On the contrary, memory-based models (Rossi et al., 2020) update node representations sequentially following the event stream, thus is significantly faster than graph-based methods yet might lose important graph information. Motivated by these observations, we propose an expressive but efficient model, Temporal Neural Common Neighbor (TNCN), to combine the merits of both. By integrating a memory-based backbone with Neural Common Neighbor (Wang et al., 2023), TNCN learns expressive pairwise representations while maintaining high efficiency akin to memory-based models. Consequently, TNCN is suitable for large-scale temporal graph link prediction.
We conducted experiments on five large-scale real-world temporal graph datasets from TGB. TNCN achieved new state-of-the-art results on three datasets and overall ranked first compared to 10 competitive baselines, demonstrating its effectiveness. To examine its scalability, we selected datasets with temporal edges ranging from $\mathcal{O}(10^{5})$ to $\mathcal{O}(10^{7})$ and node numbers ranging from thousands to millions. We found that TNCN achieved 1.9$\sim$4.7x speedup in training and 2.1$\sim$6.4x speedup in inference compared to graph-based models, while maintaining a similar scale of time consumption as memory-based models.
2 Related Work
2.1 Memory-based Temporal Graph Representation Learning
Temporal graph learning has garnered significant attention in recent years. A classic approach in this domain involves learning node memory using continuous events with non-decreasing timestamps. Kumar et al. (2019) propose a coupled recurrent neural network model named JODIE that learns the embedding trajectories of users and items. Another contemporary work DyRep (Trivedi et al., 2019) aims to efficiently produce low-dimensional node embeddings to capture the communication and association in dynamic graphs. Rossi et al. (2020) introduces a memory-based temporal neural network known as TGN, which incorporates a memory module to store temporal node representations updated with messages generated from the given event stream. Apan (Wang et al., 2021b) advances the methodology by integrating asynchronous propagation techniques, markedly increasing the efficiency of handling large-scale graph queries. EDGE (Chen et al., 2021) emerges as a computational framework focusing on increasing the parallelizability by dividing some intermediate nodes in long streams each into two independent nodes while adding back their dependency by training loss. Chen et al. (2023) extend the update method for the node memory module, introducing an additional hidden state to record previous changes in neighbors. Complementing these efforts, additional contributions such as Edgebank (Poursafaei et al., 2022) and DistTGL (Zhou et al., 2023) have been directed towards formalizing and accelerating memory-based temporal graph learning methods.
2.2 Graph-based Temporal Graph Representation Learning
Subsequent works have incorporated the temporal neighborhood structure into temporal graph learning. CAWN (Wang et al., 2021c) employs random anonymous walks to model the neighborhood structure. TCL (Wang et al., 2021a) samples a temporal dependency interaction graph that contains a sequence of temporally cascaded chronological interactions. TGAT (Xu et al., 2020) considers the temporal neighborhood and feeds the features into a temporal graph attention layer utilizing a masked self-attention mechanism. NAT (Luo and Li, 2022) constructs a multi-hop neighboring node dictionary to extract joint neighborhood features and uses a recurrent neural network (RNN) to recursively update the central node’s embedding. This information is then processed by a neural network-based encoder to predict the target link. DyGFormer (Yu et al., 2023), instead, leverages one-hop neighbor embeddings and the co-occurrence of neighbors to generate features, which are well-patched and subsequently fed into a Transformer (Vaswani et al., 2017) decoder to obtain the final prediction.
2.3 Link Prediction Methods
Link prediction is a fundamental task in graph analysis, aiming to determine the likelihood of a connection between two nodes. Early investigations posited that nodes with greater similarity tend to be connected, which led to a series of heuristic algorithms such as Common Neighbors, Katz Index, and PageRank (Newman, 2001; Katz, 1953; Page et al., 1999). With the advent of GNNs, numerous methods have attempted to utilize vanilla GNNs for enhancing link prediction, revealing sub-optimal performance due to the inability to capture important pair-wise patterns such as common neighbors (Zhang and Chen, 2018; Zhang et al., 2021; Liang et al., 2022). Subsequent research has focused on infusing various forms of inductive biases to retrieve intricate pair-wise relationships. For instance, SEAL (Zhang and Chen, 2018), Neo-GNN (Yun et al., 2021), and NCN (Wang et al., 2023) have integrated neighbor-overlapping information into their design. BUDDY (Chamberlain et al., 2022) and NBFNet (Zhu et al., 2021) have concentrated on extracting higher-order structural information. Additionally, Mao et al. (2023); Li et al. (2024) have contributed to a more unified framework encompassing different heuristics.
@@ 3 Preliminaries
(Temporal Graph) Temporal graph can be typically categorized into two kinds, discrete-time (DTDG) and continuous-time (CTDG) dynamic graph. While DTDG can be represented as a special sequence of graph snapshots in the form of $\mathcal{G}=\{(\mathcal{G}_{1},t_{1}),(\mathcal{G}_{2},t_{2}),\cdots,(\mathcal{G}_{N},t_{N})\}$, it can always be transformed into its corresponding form in CTDG. So here we mainly focus on continuous time temporal graph. We usually represent a CTDG as a sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),(u_{2},v_{2},t_{2}),\cdots,(u_{n},v_{n},t_{n})\}$, where $u,v$ stand for source and destination nodes and $\{t_{i}\}$ are chronologically non-decreasing. We use $\mathcal{V}$ to denote the entire node set, and $\mathcal{E}$ the entire edge set. Note that each node or edge can be attributed, that is, there may be node feature $x_{u}$ for $u$ or edge feature $e_{u,v}^{t}$ attached to the event $(u,v,t)$.
@@
(Temporal Neighborhood) Given the center node $u$, the $k$-hop temporal neighbor set $(k\geq 0)$ before time $t$ is defined as $N_{k}^{t}(u)=\{v\ |\ \text{SPD}^{t}(u,v)=k\}$, where $\text{SPD}^{t}(u,v)$ refers to the shortest path distance using edges before time $t$. Then the $k$-hop common neighbors between node $u$ and $v$ are defined as $\text{CN}_{k}^{t}(u,v)=\mathop{\cup}\limits_{2(k-1)<k_{1}+k_{2}\leq 2k,k_{1},k_{2}\leq k}(N_{k_{1}}^{t}(u)\cap N_{k_{2}}^{t}(v))$, i.e., all nodes on at least one $2k$-distance path between $u,v$. Note that for $k=0$, we define the 00-hop common neighbor of $u$ and $v$ as $\text{CN}_{0}^{t}(u,v)=\{u,v\}$ if there exists at least one edge between $u$ and $v$ before time $t$, and $\emptyset$ otherwise. Finally, the $K$-hop temporal neighborhood is defined as $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$.
@@ 4 Methodology
In this section, we introduce our Temporal Neural Common Neighbor (TNCN) model. TNCN comprises several key modules: the Memory Module, the Temporal CN Extractor, and the NCN-based Prediction Head. Special attention is given to the Temporal CN Extractor, designed to efficiently extract temporal neighboring structures and obtain multi-hop common neighbor information. The pipeline of TNCN is illustrated in Figure 2.
Figure 2: Pipeline of TNCN. TNCN operates through a sequential update and prediction framework that processes successive batches of messages. During the update phase, TNCN updates the neighbor dictionary and the node memory representations. In the prediction phase, the model retrieves neighbors to identify common neighbors, thereafter leveraging the representations of the target nodes and their common neighbors for prediction.
@@ 4.1 Memory Module
The memory module stores node memory representations up to time $t$. When a new event occurs, the memory of the source and destination nodes is updated with the message produced by the event. The computation can generally be represented by the following formulas:
$msg_{src}^{t}(u,v)=\textit{msgfunc}_{src}(e^{t}_{u,v}),$ $msg_{dst}^{t}(u,v)=\textit{msgfunc}_{dst}(e^{t}_{u,v}).$ (1)
$mem^{t}_{u}=\textit{upd}_{src}(mem^{t-}_{u},msg_{src}^{t}(u,v)),$ $mem^{t}_{v}=\textit{upd}_{dst}(mem^{t-}_{v},msg_{dst}^{t}(u,v)),$
where $mem^{t-}_{u}$ stands for the embedding of node $u$ before time $t$. Here, msgfunc is a learnable function, such as a linear projection or a simple identity function. Note that for different edge directions, i.e., from source to destination and vice versa, the msgfunc and updfunc can be learned separately. The memory module aids the model in managing both long-term and short-term dependencies, thereby reducing the likelihood of forgetting. During training and inference, node memory evolves dynamically as events occur. Updates to this module reflect the dynamic nature of the temporal graph.
4.2 Temporal CN Extractor
@@
Our Temporal CN Extractor can efficiently perform multi-hop common neighbor extraction and aggregate their neural embeddings.
Extended Common Neighbor.
@@
The definition of multi-hop common neighbors (CN) is given in Definition 3.3, extending the traditional 1-hop CN (i.e., nodes on 2-paths between $u$ and $v$) to $k$-hop CN (nodes on $2k$-paths between $u$ and $v$). Additionally, we define zero-hop CN to be $u$ and $v$ themselves, i.e., if there is at least one previous event between node $u$ and node $v$, then $u$ and $v$ are considered each other’s 0-hop common neighbors. The 0-hop CN not only records the historical interactions between two nodes, but also reveals the frequency of their interactions.
Efficient CN Extractor.
@@
The CN Extractor is a crucial component of the TNCN model, contributing significantly to its high performance. It can efficiently gather pertinent information about a given center node and extract multi-hop common neighbors for a source-destination pair.
@@
For each relevant node $u$, the extractor stores its historical interactions with other nodes as both source and destination. After a batch of events is processed by the model, the storage is updated with the latest interactions. This allows us to maintain a record of all historical interactions up to a certain timestamp, effectively constructing a dynamic lookup dictionary for fast retrieval during subsequent inference. To conserve memory, we can opt to save only the most recent events and relevant nodes.
During the inference stage, TNCN extracts temporal neighborhood information, i.e., historical interactions related to the query node batch. This information is organized as a Sparse Tensor, on which we perform self-multiplication to generate high-order adjacency connectivity. We use sparse tensor operations to obtain multi-hop common neighbors (CNs) and their embeddings, separating CNs from different hops. To extract the 0-hop common neighbors mentioned above in a unified way, we introduce an additional self-loop for each related node. This operation allows us to use the same tensor multiplications to obtain 0-hop CNs.
@@
The utilization of Multi-hop Common Neighbors significantly boosts TNCN’s performance, resulting in higher scores in temporal link prediction tasks. Furthermore, by employing sparse tensors, our model achieves substantial reductions in both storage requirements and computational complexity, thereby decreasing time consumption and enhancing efficiency.

### v2

@@ Abstract
Temporal graphs are widespread in real-world applications such as social networks, as well as trade and transportation networks. Predicting dynamic links within these evolving graphs is a key problem. Many memory-based methods use temporal interaction histories to generate node embeddings, which are then combined to predict links. However, these approaches primarily focus on individual node representations, often overlooking the inherently pairwise nature of link prediction. While some recent methods attempt to capture pairwise features, they tend to be limited by high computational complexity arising from repeated embedding calculations, making them unsuitable for large-scale datasets like the Temporal Graph Benchmark (TGB). To address the critical need for models that combine strong expressive power with high computational efficiency for link prediction on large temporal graphs, we propose Temporal Neural Common Neighbor (TNCN). Our model achieves this balance by adapting the powerful pairwise modeling principles of Neural Common Neighbor (NCN) to an efficient temporal architecture. TNCN improves upon NCN by efficiently preserving and updating temporal neighbor dictionaries for each node and by using multi-hop common neighbors to learn more expressive pairwise representations. TNCN achieves new state-of-the-art performance on Review from five large-scale real-world TGB datasets, 6 out of 7 datasets in the transductive setting and 3 out of 7 in the inductive setting on small- to medium-scale datasets. Additionally, TNCN demonstrates excellent scalability, outperforming prominent GNN baselines by up to 30.3 times in speed on large datasets. Our code is available at https: //github.com/GraphPKU/TNCN.
@@ 1 Introduction
Temporal graphs are increasingly employed in contemporary real-world applications like social and transaction networks, which evolve dynamically and exhibit distinct characteristics over time. Concurrently, Graph Neural Networks (GNNs) [scarselli2008graph] have emerged as prominent tools for graph representation learning, typically learning node embeddings by iteratively aggregating information from neighbors and demonstrating strong performance on various tasks. However, the temporal dimension introduces significant challenges; the discrete or continuous timestamps associated with graph edges define an evolutionary process and impose causality constraints, rendering many static GNN methodologies not directly applicable. This has led to the development of specialized temporal GNNs. Consequently, specialized temporal GNNs have been developed, often focusing on achieving both effective representation learning and computational efficiency for these large, dynamic structures. For instance, memory-based approaches like Temporal Graph Networks (TGN) [kumar2019predicting, trivedi2019dyrep, rossi2020temporal] are designed to efficiently learn short- and long-term dependencies, while Transformer-based models [wang2021tcl, xu2020inductive] utilize attention mechanisms to capture complex relationships.
@@
Despite their advancements, many specialized temporal GNNs primarily generate node-wise representations. While effective for certain tasks, this focus can be insufficient for link prediction, where accurately capturing the relationship between pairs of nodes is critical. Node-wise methods may fail to distinguish between nodes that have similar individual features or local neighborhoods but different propensities to connect with a target node (as illustrated in Figure 1), thereby overlooking crucial relational patterns. Acknowledging such limitations, approaches from static graph link prediction, like the labeling trick [zhang2018link, zhang2021labeling] that emphasizes pair-wise representations, have shown considerable success. Consequently, researchers have extended these pair-wise learning concepts to temporal graphs, often by leveraging information from the evolving local neighborhoods of node pairs [wang2021inductive, luo2022neighborhood, yu2023towards]. However, these more expressive graph-based temporal models frequently incur substantial computational and memory costs, arising from the need to extract and process temporal neighborhood information for each prediction, which can hinder their application to large-scale scenarios.
This trade-off between computationally demanding expressive models and more efficient but potentially less informative ones highlights a critical need. Addressing this, we propose the Temporal Neural Common Neighbor (TNCN) model, which is designed to achieve both high efficiency and strong expressive power for temporal link prediction. TNCN builds upon a memory-based backbone, ensuring operational efficiency comparable to sequential update models. Crucially, it incorporates a Neural Common Neighbor component [wang2023neural]. This component is augmented with advanced operational techniques and extended for multi-hop common neighbor consideration, allowing TNCN to effectively model sophisticated link heuristics and learn detailed pairwise representations while retaining the efficiency of its memory-based foundation. As a result, TNCN is well-suited for large-scale temporal graph link prediction.
[…]
@@ 2 Preliminaries
(Temporal Graph) We mainly focus on the continuous time dynamic graph (CTDG). A CTDG can be typically represented as a sequence of interaction events: $\mathcal{G}=\{(u_{1},v_{1},t_{1}),\cdots,\\ (u_{n},v_{n},t_{n})\}$, where $u,v$ stand for source and destination nodes and $\{t_{i}\}$ are chronologically non-decreasing timestamps. Note that each node or edge can be attributed, that is, there may be node feature $x_{u}$ for $u$ or edge feature $e_{u,v}^{t}$ attached to the event $(u,v,t)$.
@@
(Temporal Neighborhood) Given the center node $u$, the $k$-hop temporal neighbor set $(k\geq 0)$ before time $t$ is defined as $N_{k}^{t}(u)$. A node $v$ is in $N_{k}^{t}(u)$ if there exists a $k$-length path between $u$ and $v$, i.e. $\exists(u,w_{1},w_{2},\cdots,w_{k-1},v)$ where $w_{i}\neq w_{j},\forall i\neq j$. We also define the $(\mathbf{i},\mathbf{j})$-hop common neighbor set as follows: $w$ is an $(i,j)$-hop temporal common neighbor of $u$ and $v$ at time $t$ if $w\in N_{i}^{t}(u)$ and $w\in N_{j}^{t}(v)$. For simplicity we will denote the set as $\text{CN}_{(i,j)}^{t}(u,v)=N_{i}^{t}(u)\cap N_{j}^{t}(v)$. Note that for $i=0$ (or $j=0$ similarly), we define the 0-hop temporal neighbor set as $N_{0}^{t}(u)=\{u\}$, and the $(0,j)$-hop common neighbor of $u$ and $v$ as $\text{CN}_{(0,j)}^{t}(u,v)=N_{0}^{t}(u)\cap N_{j}^{t}(v)=\{u\}\cap N_{j}^{t}(v)$. Finally, the $K$-hop temporal neighborhood of node $u$ at time $t$ is defined as: $\mathop{\cup}\limits_{k=0}^{K}N_{k}^{t}(u)$.
With $(i,j)$-hop neighborhood information, we can perceive the local structure to a large extent and distinguish the difference between multi-hop common neighbors more precisely.
Definition 2.4.
Memory-based Backbone. Memory-based backbone has been widely adopted by various methods like [kumar2019predicting, rossi2020temporal] to tackle dynamic graph learning. Its core component is the memory module that stores the node memory representations up to a certain time $t$. When a new event occurs, the memory of the source and destination nodes is updated with the message produced by the event. The computation can generally be represented as follows:
$msg_{src}^{t}(u,v)=\textit{msgfunc}_{src}(e^{t}_{u,v}),$ $mem^{t}_{u}=\textit{upd}_{src}(mem^{t-}_{u},msg_{src}^{t}(u,v));$ (1)
$msg_{dst}^{t}(u,v)=\textit{msgfunc}_{dst}(e^{t}_{u,v}),$ $mem^{t}_{v}=\textit{upd}_{dst}(mem^{t-}_{v},msg_{dst}^{t}(u,v)).$
where $msg^{t}$ stands for the message of the event, $mem^{t-}_{u}$ for the embedding of node $u$ before time $t$.
@@ 3 Methodology
Now we introduce our Temporal Neural Common Neighbor model. TNCN comprises several key modules: the classic Memory Module, the Temporal NCN Module with efficient CN Extractor, and the NCN-based Prediction Head. Special attention will be given to the Temporal CN Extractor, which is designed to efficiently extract temporal neighboring structures and obtain multi-hop CN information. The pipeline is illustrated in Figure 2. A pseudocode is attached in Appendix F.
Figure 2: Pipeline of TNCN. TNCN operates through a sequential update and prediction framework that processes successive batches of messages. During the update phase, TNCN updates the neighbor dictionary and the node memory. In the prediction phase, the model retrieves neighbors to identify common neighbors, leveraging the representations of the target nodes and their CNs for prediction.
@@ 3.1 Memory Module
Different from the static GNN used in traditional NCN, our model TNCN adopts a memory-based backbone from [rossi2020temporal] to efficiently store and update the node memory, eliminating the need for repeated computation of node embeddings within successive temporal batches.
Conforming to the standard pipeline in section 2.4 of processing the node memory, for later link prediction or other downstream tasks, node embeddings can be obtained from their memory:
$emb^{t}_{u}=NN(mem^{t-}_{u},\mathop{\cup}\limits_{v\in N^{t}_{1}(u)}mem^{t-}_{v},\mathop{\cup}\limits_{t^{\prime}<t}[e^{t^{\prime}}_{u,v}\ ||\ \phi(t-t^{\prime})]),$ (2)
where $||$ stands for the concatenation operation. Here NN has multiple choices, like Identity or simple static GNN [defferrard2016convolutional, velickovic2017graph, hamilton2017inductive]. In our implementation, we adopt Graph Transformer Convolution [shi2020masked], which can pay more attention to the relation between different nodes. And we choose the time encoding $\phi$ presented in Time2Vec [kazemi2019time2vec] as TGN does.
3.2 Temporal NCN Module with Efficient CN Extractor
@@
Our Temporal NCN Module can efficiently perform multi-hop common neighbor extraction with its CN extractor, and aggregate their neural embeddings to attain the node features.
@@
Extended Common Neighbor. The definition of multi-hop common neighbors (CN) is given in Definition 2.3, extending the traditional (1,1)-hop CN (i.e., nodes on 2-paths between $u$ and $v$) to arbitrary $(i,j)$-hop CN. Additionally, we define the zero-hop neighbor of a central node, i.e., $u$ is considered as a neighbor of itself, which will be utilized to calculate CNs with other nodes. Given source node $u$ and target node $v$, the (0,1)-hop and (1,0)-hop CN not only records the historical interactions between two nodes, but also reveals the frequency of their interactions.
@@
Efficient CN Extractor. The CN Extractor is a crucial component of the TNCN model, contributing significantly to its high performance and scalability. It can efficiently gather pertinent information about a given center node and extract multi-hop common neighbors for a source-destination pair.
@@
For each relevant node $u$, the extractor stores its historical interactions with other nodes as both source and destination. After a batch of events is processed by the model, the storage is updated with the latest interactions. This allows us to maintain a record of all historical interactions up to a certain timestamp, effectively constructing a dynamic lookup dictionary for fast retrieval during subsequent inference. To strike a balance between memory consumption and model capacity, we save only the most recent $K$ events and relevant nodes for each center node, where $K$ is a hyperparameter determined by the specific dataset.
To implement an efficient batch CN extractor, we organize the historical interactions in a Sparse Tensor, representing the temporal adjacency matrix. Then we perform self-multiplication to generate high-order adjacency connectivity. Sparse matrix hadamard product is finally employed to obtain separate $(i,j)$-hop CNs. All these operations can be efficiently implemented by sparse tensor operators and are supported by GPU to facilitate fast, batch processing. Now we show the detailed procedure as follows. For some special and higher-order cases analysis, please refer to Appendix G.
Details of Common Neighbor Extraction. Our temporal CN extractor begins with a sparse matrix $A$ constructed from the interactions of related nodes. We then include three stages to precisely generate arbitrary $(i,j)$-hop CNs:
1. Generate up to $k$-hop neighbors. The original matrix $A$ only includes 1-hop neighbors. To extend this, we: (a) Use self-loops for 0-hop neighbors, denoted as $A^{0}$. (b) Perform sparse matrix multiplication (i.e., $A^{k}$) to include arbitrary k-hop neighbors. Combining them, we obtain an updated neighborhood matrix set $\hat{A}=\{A^{i}\}_{i=0}^{K}$.
2. Extract neighbors for each source and destination node with corresponding indices in the same batch. Assume that we require the $k$-th hop neighbors of node $u$, then vector $A^{k}[id(u)]$ is the result, where $id(u)$ stands for the reindexed id for node $u$. $A^{k}[id(u)][id(v)]=w>0$ if $v$ is a $k$-hop neighbor of $u$, otherwise this element is 0. $w$ represents the historical interaction frequency.
3. Obtain arbitrary $(i,j)$-hop CNs. We can perform hadamard product of $A^{i}[id(u)]$ and $A^{j}[id(v)]$ to acquire different hops of CNs. The operator can extract corresponding CNs for source-destination node pairs in a batch parallelly.
By re-indexing the node IDs when generating $\hat{A}$ to prevent conflicts, the CN extractor can conduct the sparse matrix calculation, which are all performed in a Torch style that supports batch operations, thus enhancing parallelism and efficiency.
@@
The utilization of Multi-hop Common Neighbors significantly boosts TNCN’s performance, resulting in higher scores in temporal link prediction tasks. Furthermore, by employing sparse tensors, our model achieves substantial reductions in both storage requirements and computational complexity, thereby decreasing time consumption and enhancing efficiency. We also give a comparison between our TNCN and traditional NCN in Table 12 in Appendix D.4.